OpenStudy (anonymous):

True or False? The function f(x)=a^2 -4 will never cross the x-axis.

OpenStudy (anonymous):

there is no \(x\) in \(a^2-4\)

OpenStudy (anonymous):

oops! Sorry, I meant a^x

OpenStudy (anonymous):

\(f(x)=2^x-4\) would cross the \(x\) axis if \(x=2\) since \(2^2-4=0\)

OpenStudy (anonymous):

Then it is false, correct?

OpenStudy (anonymous):

depends if there is any contsraints with the questions

OpenStudy (anonymous):

constraints*

OpenStudy (anonymous):

None that were mentions. All I was given was the statement: The function f(x)=a^x-4.

OpenStudy (anonymous):

well then false would the the ideal choice

OpenStudy (anonymous):

Thank you both